OpenStudy (anonymous):

what is the slope of the line passing through the points (2,7) and (1,3) a.2/7 b.3/4 c.4/3 d.1/3

OpenStudy (anonymous):

I think its 2/7 help?

OpenStudy (anonymous):

the slope is 4 it is f because you put the rise over the run so you have to go up 4 and over 1 that means it is a slope of 4

OpenStudy (anonymous):

so would that mean that it is 4/3?

OpenStudy (anonymous):

no it would be 4/1

OpenStudy (anonymous):

oh yeah I get it I forgot to do that. you mind helping me with a couple more please?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

http://prntscr.com/4z9v57 a.y= 3/2x + 3 b.y=-3/2x+ 3 c.y=2/3x + 3 d.=-2/3x + 3

OpenStudy (anonymous):

what are you asking

OpenStudy (anonymous):

is the url supposed to be a pic

OpenStudy (anonymous):

what is the correct equation of the line shown below

OpenStudy (anonymous):

yes the url is the diagram

OpenStudy (anonymous):

any ideas? I think its c to be honest but im horrible at geometry so im most likely wrong...